sammixboo (sammixboo):

The temperature in a freezer is falling steadily at a rate of two degrees Fahrenheit per minute. The function y equals twenty-five minus two x gives the temperature y after x minutes. Which set of terms best describes the function y equals twenty-five minus two x in this scenario? discrete and linear discrete and non-linear continuous and linear continuous and non-linear

sammixboo (sammixboo):

Somebody please help...

sammixboo (sammixboo):

@mathmale

sammixboo (sammixboo):

@phi

OpenStudy (phi):

y equals twenty-five minus two x gives the temperature y after x minutes. It is easier to read if you write it like this y = 25 - 2x you can also write it y = -2 x + 25 does that equation remind you of y= mx + b ???

sammixboo (sammixboo):

Yes it does.

OpenStudy (phi):

what kind of equation is it ? (Please say, it is the equation of a line)

OpenStudy (phi):

which means LINEAR discrete means the answers are specific numbers, like 1,2,3, with gaps continuous means y can be any number (continuous... no gaps... like a line)

sammixboo (sammixboo):

I got it thanks

sammixboo (sammixboo):

continuous and linear

OpenStudy (phi):

yes